OpenStudy (anonymous):

Differentiate and evaluate at x=0, y=(x^2-2x+9) e^x?

OpenStudy (anonymous):

the derivatice should be \[\huge (x^2-2x+9)*e^x+(2x-2)\]

OpenStudy (anonymous):

or

OpenStudy (anonymous):

\[\huge (x^2+7)*e^x\]

OpenStudy (anonymous):

I thought it should be \[e^x (x^2-2x+9)+e^x (2x-2)=e^x(x^2-7)\]

OpenStudy (anonymous):

if x=0 then it will be

OpenStudy (anonymous):

You don't do anything with the e^x?

OpenStudy (anonymous):

Yeah, you don't.

OpenStudy (anonymous):

\[\huge 7e^x\]

OpenStudy (anonymous):

You can try that out from first principle.

OpenStudy (anonymous):

apply the product rule first

OpenStudy (anonymous):

Whats the product rule?

OpenStudy (anonymous):

I can't find it

OpenStudy (anonymous):

\[\frac{d}{dx}f(x)g(x)=f'(x)g(x)+f(x)g'(x)\]

OpenStudy (anonymous):

same as quotient but without the denominator?

OpenStudy (anonymous):

d(uv)=u(dv)+v(du)

OpenStudy (anonymous):

Yeah, you can derive the quotient rule from product rule.

OpenStudy (anonymous):

They are the same literally.

OpenStudy (anonymous):

where did you guys get (2x-2)?

OpenStudy (anonymous):

\[\frac{d}{dx}(x^2-2x+9)=2x-2\]

OpenStudy (anonymous):

yup!

OpenStudy (anonymous):

Oh wow, lol hahah I see now, thanks

OpenStudy (anonymous):

Does it matter where the e^x is? like in front of the parenthesis or behind